H2Cre.co HzCvcoph3P ....phpPn3 PhP on- H HD. CH Co(73-C,Hz)(CO),HIC H Hac HOOC COOH Note: Divalent NR2 is an X ligandUse both the ionic and covalent counting systems as was done in the examples in class to give electron counts for the metal and each type of ligand. Use these to determine the total electron count for these organometallic compounds (watch for oxidation state of cationic complexes!  Also, not all are 18 e‐ ):
Answer #1

- Q Br zo = 4 - -BY 2Br = 2 Covalent method ionic methol 12+ = 2Br = 4 2Cp = 10 200 = 12 total ě - 16 - To - - count 1 Fe2+ =CH3 Fe (0) = 8 Fe=8 4 N = 8 / i 40 = 8 1 16 HC, totale = 16 count Hooc COOH Pd = 10 of= 3 I pd4* = 6 1 BE = -a- 74 totallé 14